An interior decorator is buying wallpaper for two square walls. One wall is 4 feet wider than the other. If he/she needs 170 square feet of wall paper, what are the dimensions of the wall?Show solution
Answer by Boreal(15235)

smaller wall has side x
larger wall has side x+4
area of smaller wall is x^2; area of larger wall is (x+4)^2=x^2+8x+16.
These areas add to 170
So 2x^2+8x+16=170
x^2+4x+8=85
x^2+4x-77=0
(x+11)(X-7)=0
x=7 feet, only positive root.
one wall is 7 x 7 feet or 49 ft^2
other wall is 11 x 11 feet or 121 ft^2
They add to 170 ft^2.